OpenStudy (anonymous):

A man on the fifth floor of a building shouts down to a person on the street. If the angle of elevation from the street to the man in the building is 35° and the man in the building is 40 feet up, about how far away from the building is the person on the street?

OpenStudy (anonymous):

A) 28 feet B) 55 feet C) 56 feet D) 57 feet

OpenStudy (anonymous):



OpenStudy (anonymous):

\[\sin(35)=\frac{40}{x}\iff x=\frac{40}{\sin(35)}\] and a calculator

OpenStudy (anonymous):

...i can't get it.

OpenStudy (anonymous):

hmmm i don't get any of those answers...

OpenStudy (anonymous):

are you sure those are the right choices?

OpenStudy (anonymous):

Here's the picture that is with the question.

jimthompson5910 (jim_thompson5910):

you use tangent, not sine

jimthompson5910 (jim_thompson5910):

tan(35) = 40/x x*tan(35) = 40 x = 40/tan(35) x = ???

OpenStudy (anonymous):

D ?

OpenStudy (anonymous):

oooh how far is the building from the man, not the two man doh

jimthompson5910 (jim_thompson5910):

yeah somewhat of a trick question

OpenStudy (anonymous):

what @jim_thompson5910 said i should learn to read

jimthompson5910 (jim_thompson5910):

57 ft is correct

OpenStudy (anonymous):

@satellite73 thank you for trying to help! @jim_thompson5910 Thank you for giving the answer!

OpenStudy (anonymous):

\[x=\frac{40}{\tan(35)}\] http://www.wolframalpha.com/input/?i=40%2Ftan%2835%29